OpenStudy (anonymous):

Janice deposited $750 in a savings account that earns 3.5% simple interest. How much interest has Janice earned by the end of the first year?

OpenStudy (gschibby):

\[750*1,035=?\]

OpenStudy (anonymous):

How did you figure this? your right i just would like to get it to. lol

OpenStudy (gschibby):

Well, 3,5% is the same as 0,035, but since you already have the deposit, you have to have to 1, too. You could also write it out like this \[Interest = 750*0,035=262,5\] that would be the correct answer to your question, but the amount on the account would be:\[Amount = savings+interest = 750+(750*0,035)\] It's just easier to write\[750*1,035\] directly ;P

OpenStudy (anonymous):

How would i do this one? Aaron deposited $591.07 in a savings account that earns 3.7% simple intrest. What is Aaron's account balance after seven years.

OpenStudy (gschibby):

Hm, I'm a little sketchy on what "simple interest" means :P

OpenStudy (gschibby):

If it means that the interest won't change in seven years it's easy:\[Balance = 591,07*1,037^{7}\]

OpenStudy (gschibby):

wich is 762,24 ;)

OpenStudy (anonymous):

Susan currently has an account balance of 1336.60. Her initial deposit on the account was $820 and it earned 4.5% simple interest. How long has she held the account?

OpenStudy (gschibby):

Since the question is still open, I'll just answer it :P Set it up like this: \[1336.60=820*1.045^{x}\] Then you solve for x: \[1.048^{x}=\frac{ 1336.6 }{ 820 }=1.63\] \[\ln 1.045^{x}=\ln1.63\] That way, when using natural logarisms, you can move the x down: \[x*\ln 1.045=\ln 1.63\] \[x=\frac{ \ln 1.63 }{ \ln 1.045 }=11.0998\approx11.1\] This means that Susan has held the account for 11.1 years :)

OpenStudy (gschibby):

Then you can check it: \[820*1.045^{11.1}=1336.6\] ;)